# Disentangling Rich Dynamics from Feature Learning: A Framework for Independent Measurements

### v2

@@ 3.1 Low rank bias as richness measure
See Appendix D for proofs and further discussions.
@@
Although the ideal criteria of $\mathcal{T}$ is $\mathcal{T}_{MP}$ implies neural collapse criteria as a special case, they differ in general: we measure how well features (random variables or functions) express the learned function, while neural collapse concerns how training feature vectors represent class mean vectors. Based on function space, our measure extends to test data, enables feature quality assessment (Eq. 5), and is more robust (Fig. 2), extending beyond neural collapse (see Appendix D).
@@ 3.2 Visualization through decomposed features
So far, we have focused on a metric for dynamical richness. While metrics offer quantitative summaries, visualization by components reveals underlying structure and trends obscured by raw number — helping interpret whether an observed metric value is meaningfully small or large enough (e.g., Fig. 4). We also aim to incorporate feature improvement (thus quality) into these visualizations to encourage future work on its relationship with dynamical richness.
@@
Bengio et al. [1] described feature learning as the process of learning better representations for a downstream model, such as the classifier. Building on this idea, we view the earlier layers, represented by the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$, as providing improved features for the final linear layer. Combined with our $\mathcal{T}$-dependent richness metric, this motivates a visualization method that extends the linear model analysis of $\mathcal{T}$ (Appendix C).
@@
Our framework quantifies three aspects (cumulative quality $\Pi^{*}(k)$, cumulative utilization $\hat{\Pi}(k)$, and relative eigenvalue $\rho_{k}/\rho_{1}$) of the last-layer features:
$(i):\Pi^{*}(k)=\sum_{j=1}^{k}\frac{\braket{e_{k}}{P_{\mathcal{H}^{*}}[e_{k}]}^{2}}{\textrm{dim}(\mathcal{H}^{*})},\quad\quad(ii):\hat{\Pi}(k)=\sum_{j=1}^{k}\frac{\braket{e_{k}}{P_{\hat{\mathcal{H}}}[e_{k}]}^{2}}{\textrm{dim}(\hat{\mathcal{H}})},\quad\quad(iii):\rho_{k}/\rho_{1},$ (5)
@@
where $\mathcal{H}^{*}=span\{f^{*}_{1},\ldots,f^{*}_{C}\}$ and ^$\hat{\mathcal{H}}=span\{\hat{f}_{1},\ldots,\hat{f}_{C}\}$ are the target and learned function spaces, and $P_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{H}$ is the projection operator onto the space $\mathcal{H}$. All cumulative measures lie in $[0,1]$, reflecting how well the top $k$ features span the respective space. Notably, $\Pi^{*}(k)$ is analogous to the task-model alignment, quantifying the contribution of the top features in expressing the target.
@@
These three measures together capture complementary views of feature learning. The cumulative quality ($\Pi^{*}(k)$) reflects how well the features align with the task (i.e., task-model alignment). The utilization ($\hat{\Pi}(k)$) indicates how many features are used by the final layer, while the relative eigenvalues ($\rho_{k}/\rho_{1}$) show their relative magnitudes or importance. The latter two ($\hat{\Pi}$ and $\rho_{k}/\rho_{1}$) decompose the deviation from $\mathcal{D}_{LR}=0$ condition by visualizing how many features are used and how many are significant (non-negligible).
3.3 Approximating the metrics in practice
@@
To calculate the eigenvalues and eigenfunctions (and thus the quality, utilization, and intensity) for a feature map $\Phi$, we need the true input distribution $q$, which is, in principle, inaccessible. However, we can use Nyström method [37, 38] for a sufficiently large sample size of $n>p$ to approximate the eigenfunctions and eigenvalues of interest. We can create a $p\times p$ empirical self-covariance matrix $\Sigma$ from $n$ samples of $\Phi(x)$. By diagonalizing $\Sigma=USU^{T}$, we can obtain the empirical eigenvectors and eigenvalues, which can be used to approximate the true eigenvalues and eigenfunctions:
@@
1:  $\Phi,D_{tr}$ (Prepare a feature map $\Phi$ and $D_{tr}$: $n$ samples of training set)
2:  $\phi\leftarrow\Phi(D_{tr})$ (forward transform the input samples to feature vectors)
3:  $U,S,U^{T}\leftarrow SVD(\phi\phi^{T})$ (perform singular value decomposition)
4:  $[u_{1},u_{2},\ldots,u_{p}]\leftarrow U$ (find the column vectors of $U$)
@@
Note that eigenfunctions $e_{k}^{\prime}s$ are defined beyond the training samples used to compute $\phi$ and differ from the eigenvectors of the empirical self-covariance matrix. This enables us to evaluate feature quality — via inner products with the target function — on the test set rather than the training set, a capability not available for empirical eigenvectors.
Time complexity.
The algorithm is computationally efficient in practice. The main cost in computing Eq. 6 is the SVD, typically with time complexity $\mathcal{O}(p^{3})$. A good approximation requires only $n=\mathcal{O}(p)$ datapoints—more than the last layer width. Performing SVD on $\phi$ instead of $\phi\phi^{T}$ yields $U$ and $\sqrt{S}$ with $\mathcal{O}(p^{3})$ cost. Because last layers are typically narrow (e.g., $p\approx 10^{3}$) and the algorithm requires just $n=\mathcal{O}(p)$ forward passes, the overall computation is lightweight.
4 Comparison to prior measures of rich dynamics
Beyond performance — often used as a proxy for rich dynamics [2, 10] — three common metrics are: (1) similarity to the initial kernel, $\mathcal{S}_{init}=CKA(K_{\text{init}},K_{\text{learned}})\in[0,1]$ [39], (2) the parameter norm $\|\theta\|_{F}^{2}$ [12], and (3) class separation from neural collapse, $NC1=Tr(\Sigma_{b}^{\dagger}\Sigma_{W})$ [6, 40, 41, 42, 43], where $\Sigma_{b}$ and $\Sigma_{W}$ are the inter- and intra-class covariances (Proposition 1). Prior metrics have limitations as an independent measure of richness as they rely on the initial kernel, parameter norms, or class labels.
Epoch 0 (init) 200
Train Acc.$\uparrow$ 10.0% 10.2%
Test Acc.$\uparrow$ 9.85% 10.0%
$\bm{\mathcal{D}_{LR}}$$\downarrow$ 0.59 1.0
$\mathcal{S}_{init}$$\downarrow$ 1.0 0.20
$\|\theta\|_{F}^{2}$$\downarrow$ 3.1 $\cdot 10^{3}$ 2.2 $\cdot 10^{-5}$
$NC1$$\downarrow$ $1.2\cdot 10^{5}$ $7.5\cdot 10^{-14}$
(a) Weight decay dynamics
$\alpha$ $2\cdot 10^{-1}$ $2\cdot 10^{0}$ $2\cdot 10^{1}$
Train Acc.$\uparrow$ 100 % 100% 100%
Test Acc.$\uparrow$ 92.7% 92.4% 88.3%
$NC1$$\downarrow$ $2.3\cdot 10^{4}$ $3.2\cdot 10^{3}$ $8.1\cdot 10^{2}$
(b) Target downscaling
Figure 2: A 4-layer MLP trained on $10^{3}$ MNIST samples. (a) Training with larger weight decay $10^{-3}$ and smaller learning rate $10^{-6}$. While our metric correctly indicates the lack of richness, all other metrics mistakenly suggest a significant rich dynamics with smaller post-training values. (b) Training with target downscaling $\alpha$, where larger $\alpha$ leads to lazier dynamics. Only our measure correctly identifies the change in laziness. For a discussion on NC1, See Appendix F. (c) Our visualization methods for target downscaling, illustrating how lazier dynamics (ii) use more features and (iii) show a slower decay of eigenvalues (ignoring $\rho_{1}$ of constant function, see Appendix D).
Figure 2(a) shows an extreme case of training an MLP on MNIST with large weight decay and a negligible learning rate. Here, the dynamics are dominated by L2 weight decay, with little meaningful learning. While existing metrics often misinterpret this as rich behavior (smaller after training), our span-based measure correctly identifies the lack of dynamical richness (bigger after training). Figure 2(b,c) presents a more practical setting, where we tune laziness via target downscaling (i.e., $y\rightarrow y/\alpha$). Prior works [2, 3] show that scaling targets by a factor $\alpha$ induces lazier training where larger $\alpha$ implies greater laziness. A good metric should capture this. In Fig. 2(b), our measure aligns with $\alpha$, while all other measures misalign with laziness. Finally, Fig. 2(c) visualizes the reduced dynamical richness by showing that more features are used and are significant as $\alpha$ increases.
Kernel distance from initialization ($\bm{\mathcal{S}_{init}}$).
Kernel deviation is a common measure of feature learning [2], often computed via the change in the neural tangent kernel (NTK) [3, 44, 45]. However, empirical NTK is computationally expensive in practice, so here we observe the change in the last-layer feature kernel instead. Importantly, kernel deviation can be misleading in deep, rich-training regimes where most practical training occurs. While it detects movement away from the initial kernel, it cannot distinguish different degrees of richness once the dynamics are non-lazy, nor does it capture the specific low-rank biases characteristic of rich feature learning.
Parameter norm ($\bm{\|\theta\|_{F}^{2}}$).
The parameter norm, used in Lyu et al. [12] as a proxy for richness, may offer a limited perspective as it ignores the learning dynamics and is unbounded in scale. Indeed, networks can enter a rich regime even as their weight norm increases [22, 46]. While smaller norms do often correlate with rich dynamics in practice, this is because small weights tend to promote rich dynamics [10, 47, 13, 4, 48] — not because rich dynamics inherently produce small norms. In fact, rich dynamics depend on broader factors such as network layer imbalance [46, 22, 11, 13].
Neural collapse measure ($\bm{NC1}$).
Separation-based metrics measure how training samples deviate relative to class boundaries [6, 40, 41, 42, 43]. They are designed for the terminal phase of training and depend heavily on class labels (thus training accuracy) and especially the output scale — making them unreliable during training. Their unbounded nature further complicates interpretation. As shown in Fig. 2(a), they are overly sensitive to output scaling, falsely suggesting increased richness under weight decay.
Our measure ($\bm{\mathcal{D}_{LR}}$).
Our proposed metric evaluates the alignment between features and the learned function, achieving its optimum when they span the same space isotropically (i.e., a scaled projection operator). Our metric exploits the low-rank bias of greedy dynamics through the alignment and is normalized between $[0,1]$. Crucially, it does so without relying on class labels, accuracy, or the initial kernel, making the measure more appealing as an independent measure of richness.
5 Observations from our visualization methods
In this section, we present a series of empirical findings to demonstrate the effectiveness of our visualization framework (Eq. 5) and to highlight observations that open revenues for future theoretical studies. We revisit several known phenomena by our visualization framework, such as the emergence of rich dynamics in the absence of underlying data structure (Fig. 3) and the transition from lazy to rich regime during grokking (Fig. 4). We also report novel observations, including the role of training sample size (Fig. 5) and batch normalization (Fig. 6) in encouraging richness, and a correlation between the quality and intensity of rich dynamics during training (Fig. 7).
For completeness, additional empirical results are provided in Appendices E and D, including evidence of low-rank bias without weight decay (Fig. 14), the impact of the learning rate on dynamical richness (Fig. 15), the richer-to-lazier transition on MLP (Fig. 16), and the dependence of low-rank bias on the dimensionality of the output function rather than the number of target classes (Fig. 12).
Rich dynamics without underlying data structure.
Fig. 3 illustrates that rich dynamics can occur independently of representation enhancement or underlying data structure. We trained ResNet18 on CIFAR-10 with varying levels of label shuffling and observed that even with fully randomized labels, the model enters the rich regime. This suggests that the dynamical low-rank bias is strong enough to collapse expressive networks into minimal representations, consistent with prior observations of neural collapse under random labeling [31, 33].
Figure 4: Grokking as a transition from the lazy to rich regime. A 2-layer transformer is trained on the modular $p$ division task. The inset in the middle shows the training and test accuracies, where green and red vertical lines indicate before and after grokking (steps 200 and 3000). Our metric in square brackets shows the transition into the rich regime, while our visualization shows a clear difference in the use of features.
Figure 5: Learning curve and feature learning metric. (a): Learning curves of ResNet18 on CIFAR10. Both error (a) and loss (b) learning curves show a transition to a faster-decaying power law with additional data near $n\approx 10^{3}$. This correlates with the shift in decay of the richness measure $\mathcal{D}_{LR}$ in (c), suggesting a transition into richer dynamics. A linear model (Gaussian process) was plotted in (a,b) to highlight the transition into faster-decaying learning curve.
Grokking.
Fig. 4 confirms that grokking involves a transition from a lazy, overfitting regime to a rich, generalizing one [10, 11, 12, 13]. While prior work has analyzed the dynamics, the transition into the rich regime was often inferred from the grokking behavior of the performance. Our metric provides an independent measure that aligns and supports previous studies.
Learning curves.
In Fig. 5, the learning curve follows a power law but exhibits a shift in the exponent at a dataset size $n$, which coincides with a sudden transition into the rich regime. This suggests that a sufficiently large number of data points is critical for initiating rich dynamics — a promising observation toward better understanding feature learning dynamics.
Batch normalization.
Fig. 6 shows a novel observation that VGG16 without batch norm is in the lazy regime, while adding batch norm leads to the rich regime. While the performance benefits of batch norm are well known, our method offers a new perspective on its influence over lazy vs. rich dynamics, shedding light on why it improves performance.
Rich feature learning.
Fig. 7 reveals another novel pattern in feature learning dynamics: feature quality correlates with feature intensity during training, with larger features improving faster. While it is expected that a generalizing model in the rich regime obtains a few high-quality features after training, the correlation between quality and intensity during training has not been previously observed or studied. Interestingly, it appears to contrast with the silent alignment [49], which suggests quality (dynamically) precedes intensity — highlighting a new direction for future research.

### v3

@@ 3.2 Connection to neural collapse
See Appendix˜D for proofs and further discussions.
@@
Although the ideal criteria of $\mathcal{T}$ is $\mathcal{T}_{MP}$ implies neural collapse criteria as a special case, they differ in general: we measure how well features (random variables or functions) express the learned function, while neural collapse concerns how training feature vectors represent class mean vectors. Based on the function space, our measure extends to test data, enables feature quality assessment (Eq.˜5), and is more empirically robust (Tables˜2 and 2), extending beyond neural collapse (see Appendix˜D).
4 Experiments
Here, we share our empirical results. We first compare our metric with existing metrics of richness, then we confirm that our measure empirically tracks known lazy-to-rich transitions. We further conduct experiments on various training setups to explore the relationship between training factors, rich dynamics, and performance (better representations).
4.1 Comparison to prior measures of rich dynamics
Three alternative metrics for richness measures are: (1) similarity to the initial kernel, $\mathcal{S}_{init}=CKA(K_{\text{init}},K_{\text{learned}})\in[0,1]$ (Yang and Hu, 2021), (2) the parameter norm $\|\theta\|_{F}^{2}$ (Lyu et al., 2024), and (3) class separation from neural collapse, $NC1=\mathrm{Tr}(\Sigma_{b}^{\dagger}\Sigma_{W})$ (Papyan et al., 2020; Stevens and others, 2002; He and Su, 2023; Xu and Liu, 2023; Súkeník et al., 2024), where $\Sigma_{b}$ and $\Sigma_{W}$ are the inter- and intra-class covariances (Proposition˜1). Prior metrics depend on the initial kernel, parameter norms, or class labels, which can constrain their use as independent measures of richness.
Table 1: Richness for weight decay
Epoch 0 (init) 200
[…]
$\bm{\mathcal{D}_{LR}}\downarrow$ 0.59 1.0
$\mathcal{S}_{init}\downarrow$ 1.0 0.20
$\|\theta\|_{F}^{2}\downarrow$ 3.1 $\cdot 10^{3}$ 2.2 $\cdot 10^{-5}$
$NC1\downarrow$ $1.2\cdot 10^{5}$ $7.5\cdot 10^{-14}$
Table 2: Richness metrics for target downscaling
$\alpha$ $2\cdot 10^{-1}$ $2\cdot 10^{0}$ $2\cdot 10^{1}$
[…]
$NC1\downarrow$ $2.3\cdot 10^{4}$ $3.2\cdot 10^{3}$ $8.1\cdot 10^{2}$
Table˜2 shows an extreme case of training an MLP on MNIST with large weight decay and a negligible learning rate. Here, the dynamics are dominated by $L^{2}$ weight decay, with little meaningful learning. While existing metrics can sometimes misinterpret this as rich behavior (smaller after training), our span-based measure correctly identifies the lack of dynamical richness (bigger after training).
Table˜2 presents a more practical setting, where we tune laziness via target downscaling (i.e., $y\rightarrow y/\alpha$). Prior works (Chizat et al., 2019; Geiger et al., 2020) show that scaling targets by a factor $\alpha$ induces lazier training where larger $\alpha$ implies greater laziness. A good metric should capture this. Our measure aligns with $\alpha$, while all other measures misalign with laziness. By not relying on initial kernel, weight norm, or labels, our metric shows greater robustness in this setup. See Fig.˜11 in Appendix˜F for visualization of the dynamics in Table˜2.
Kernel distance from initialization ($\bm{\mathcal{S}_{init}}$).
Kernel deviation (Yang and Hu, 2021), albeit often the comparison of NTKs, is a common measure of rich dynamics (Chizat et al., 2019). However, it measures the deviation from the initial kernel, not how the kernel changed (e.g. weight decay alone changes the kernel in Table˜2). While theoretically appealing for near-lazy models, the metric can be less accurate in deep, rich-training regimes where most practical training occurs.
Parameter norm ($\bm{\|\theta\|_{F}^{2}}$).
Smaller parameter norms often correlate with rich dynamics in practice, similar to how the authors of Lyu et al. (2024) used it in their study of grokking. However, small weights promote rich dynamics (Kumar et al., 2024; Atanasov et al., 2024; Nam et al., 2025; Saxe et al., 2014; 2019), not the other way around. In fact, rich dynamics can occur for larger initialization (Braun et al., 2022; Dominé et al., 2025) and depends on broader factors such as layer imbalances (Kunin et al., 2024; Dominé et al., 2025; Nam et al., 2025).
Neural collapse measure ($\bm{NC1}$).
Separation-based metrics measure how training samples deviate relative to class boundaries (Papyan et al., 2020; Stevens and others, 2002; He and Su, 2023; Xu and Liu, 2023; Súkeník et al., 2024). While most similar to our metric, they are unbounded, sensitive to output scaling, and can be empirically ill-conditioned. As shown in Table˜2, these issues can lead to dramatic value shifts and hinder interpretability.
Low-rank measure ($\bm{\mathcal{D}_{LR}}$).
Our proposed metric evaluates the alignment between features and the learned function, achieving its optimum when they span the same space isotropically (i.e., a scaled projection operator). Our metric exploits the low-rank bias of rich dynamics through alignment and is normalized between $[0,1]$. Crucially, it does so without relying on class labels, accuracy, or the initial kernel, making the measure more appealing as an independent measure of richness.
4.2 Empirical findings regarding training factors
Table 3: Richness measure and performance on various setups
Task Architecture Condition Train acc. Test acc. $\mathcal{D}_{LR}$ Figure
Mod 97 2-layer transformer Step 200 (before grokking) 100% 5.2% 0.51 Fig.˜12
Step 3000 (after grokking) 100% 99.8% 0.11
CIFAR-100 ResNet18 learning rate = 0.005 100% 66.3% 0.053 Fig.˜4
learning rate = 0.05 100% 78.3% 0.025
learning rate = 0.2 100% 74.5% 0.039
CIFAR-10 ResNet18 weight decay = 0 100% 93.5% 0.05 Fig.˜13
weight decay = $10^{-4}$ 100% 94.1% 0.015
weight decay = $10^{-3}$ 100% 94.8% 0.003
CIFAR-10 ResNet18 This experiment compares architectures only. 100% 94.8% 0.026 Fig.˜14
MLP 99.8% 55.4% 0.48
CIFAR-10 ResNet18 no label shuffling 100% 95.0% 0.031 Fig.˜15
10% label shuffling 100% 66.1% 0.042
full label shuffling 100% 9.5% 0.034
MNIST CNN full backpropagation 100% 99.1% 0.043 Fig.˜16
last layer only training 99.7% 96.8% 0.51
CIFAR-100 VGG-16 without batch normalization 99.5% 21.7% 0.66 Fig.˜3
with batch normalization 100% 72.0% 0.073
Table˜3 demonstrates the practical usefulness of our performance-independent metric, summarizing the correlation between training factor, performance, and richness in various setups. The visualization of most experiments is provided in Appendix˜F.
The first row empirically confirms that our metric is a richness metric, explicitly capturing the known lazy-to-rich transition of grokking (Kumar et al., 2024; Kunin et al., 2024; Lyu et al., 2024; Nam et al., 2025).
Rows two to four confirm the assumptions that the optimal learning rate ($2^{\mathrm{nd}}$ row), the weight decay ($3^{\mathrm{rd}}$ row), and the architecture ($4^{\mathrm{th}}$ row) achieve high performance through rich dynamics (Ginsburg et al., 2018; Liu et al., 2022; He et al., 2016). Indeed, our metric makes this link explicit by having a smallest value for the optimal setting. Although such relationships are broadly recognized, they are typically only addressed implicitly.
The fifth and sixth rows, like Fig.˜1, demonstrate that rich dynamics do not strictly correlate with performance. ResNet18 on CIFAR-10 retains rich dynamics even when the labels are shuffled, and convolutional neural network (CNN) on MNIST can achieve similar performance through both lazy and rich dynamics.
The last row reveals a new observation: VGG-16 on CIFAR-100 is lazy without batch normalization but rich with it, accompanied by a significant performance gap. While the performance effect of batch normalization is empirically established, its underlying role remains debated. Our metric helps clarify this by reframing generalization in terms of the more tractable problem of rich dynamics.
Figure 2: Learning curve and feature learning metric. (a): Learning curves of ResNet18 on CIFAR-10. Both error (a) and loss (b) learning curves show a transition to a faster-decaying power law with additional data near $n\approx 10^{3}$, correlating with the shift in decay of the richness measure $\mathcal{D}_{LR}$ in (c). This agrees with theoretical study on phase transition (Rubin et al., 2024) that a sufficiently large number of data points is critical for rich dynamics — a promising observation toward better understanding feature learning dynamics. A linear model (Gaussian process) was plotted in (a,b) to highlight the transition into faster-decaying learning curve.
5 Visualization methods
While metrics provide quantitative summaries, component-wise visualization aids interpretation. For instance, how significant is the contrast between VGG-16 trained with versus without batch normalization (Fig.˜3)? Or why does a small change in $\mathcal{D}_{LR}$ correspond to large performance differences under varying learning rates (Fig.˜4)? Does the CIFAR-10 MLP begin in a lazy regime, or transition into it during training (Fig.˜14 in Appendix˜F)? To address such questions, we introduce a complementary visualization based on widely used eigendecomposition of kernel (Eq.˜1).
@@ 5.1 Visualization through decomposed features
Bengio et al. (2013) described feature learning as the process of learning better representations for a downstream model, such as the classifier. Building on this idea, we view the earlier layers, represented by the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$, as providing improved features for the final linear layer. Combined with our $\mathcal{T}$-dependent richness metric, this motivates a visualization method that extends the linear model analysis of $\mathcal{T}$. See Appendix˜C for related works and a gentle overview.
@@
Our visualization quantifies three aspects (cumulative quality $\Pi^{*}(k)$, cumulative utilization $\hat{\Pi}(k)$, and relative eigenvalue $\rho_{k}/\rho_{1}$) of the last-layer features through the eigenfunctions $e_{k}$ of $\mathcal{T}$:
$(i):\Pi^{*}(k)=\sum_{j=1}^{k}\frac{\braket{e_{k}|P_{\mathcal{H}^{*}}[e_{k}]}^{2}}{\textrm{dim}(\mathcal{H}^{*})},\quad\quad(ii):\hat{\Pi}(k)=\sum_{j=1}^{k}\frac{\braket{e_{k}|P_{\hat{\mathcal{H}}}[e_{k}]}^{2}}{\textrm{dim}(\hat{\mathcal{H}})},\quad\quad(iii):\rho_{k}/\rho_{1},$ (5)
@@
where $\mathcal{H}^{*}=\mathrm{span}\{f^{*}_{1},\ldots,f^{*}_{C}\}$ and $\hat{\mathcal{H}}=\mathrm{span}\{\hat{f}_{1},\ldots,\hat{f}_{C}\}$ are the target and learned function spaces, and $P_{\mathcal{H}}:L^{2}(\mathcal{X})\rightarrow\mathcal{H}$ is the projection operator onto the space $\mathcal{H}$. All cumulative measures lie in $[0,1]$, reflecting how well the top $k$ features span the respective space. Notably, $\Pi^{*}(k)$ is the cumulative power in Canatar and Pehlevan (2022), quantifying the contribution of the top features in expressing the target.
@@
These three measures together capture complementary views of feature learning. The cumulative quality ($\Pi^{*}(k)$) reflects how well the features align with the task. The utilization ($\hat{\Pi}(k)$) indicates how many features are used by the final layer, while the relative eigenvalues ($\rho_{k}/\rho_{1}$) show their relative magnitudes or importance. The latter two ($\hat{\Pi}$ and $\rho_{k}/\rho_{1}$) decompose the deviation from $\mathcal{D}_{LR}=0$ condition by visualizing how many features are used and how many are significant (non-negligible).
5.2 Calculating the visualization metrics
@@
To calculate the eigenvalues and eigenfunctions (and thus the quality, utilization, and intensity) for a feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$, we need the true input distribution $q$, which is inaccessible. However, we can use Nyström method (Baker and Taylor, 1979; Williams and Rasmussen, 2006) for a sufficiently large sample size of $n>p$ to approximate the eigenfunctions and eigenvalues of interest. We can create a $p\times p$ empirical self-covariance matrix $\Sigma$ from $n$ samples of $\Phi(x)$. By diagonalizing $\Sigma=USU^{T}$, we can obtain the empirical eigenvectors and eigenvalues, which can be used to approximate the true eigenvalues and eigenfunctions (Algorithm˜1):
@@
[…]
@@
Note that eigenfunctions $e_{k}^{\prime}s$ are defined beyond the training samples used to compute $\phi$ and differ from the eigenvectors of the empirical self-covariance matrix. This enables us to evaluate feature quality — via inner products with the target function — on the test set rather than the training set, a capability not available for empirical eigenvectors. For validity of approximation and dependency on sample size, see Appendix˜J. As eigendecomposition arises naturally in study of kernels, similar visualizations have appeared in prior works (e.g., (Canatar et al., 2021)); see Appendix˜G for details.
5.3 Visualization results
Interpreting visualization
In perfectly rich dynamics (e.g., blue line in Fig.˜3), $\hat{\Pi}(C)=1$ in (ii), showing that only the first $C$ features are used by the last layer. Additionally, $\rho_{2}\approx\rho_{C}$ and $\rho_{C}\gg\rho_{C+1}$ in (iii), indicating that the features have only $C$ significant dimensions. The first eigenvalue, corresponding to a constant function, is ignored because of the centering in CKA (see Appendix˜D). If the performance is high, $\Pi^{*}(C)\approx 1$ in (i), showing that first $C$ features well express the target. For the lazy regime, more eigenfunction will be used (slower capping in (ii)) and the feature occupy higher dimensions (slower decay in (iii)).